For a standard normal​ distribution, determine the following probabilities.

​a) P(z>1.41​)

​b) P(z>−0.31​)

​c)​ P(−1.81≤z≤−0.69​)

​d) P(−1.80≤z≤0.20​)
Answer #1

a) P(z > 1.41) = 1- P(z < 1.41) = 1-0.92073 = 0.07927

b) P(z > -0.3) = 1 - P(z <= -0. 3) = 1 - 0.38209 = 0.61791

c) P(−1.81 ≤ z ≤ −0.69​) = P(z < -0.69) - P(z < -1.81) = 0.24510 - 0.03515 = 0.20995

​d) P(−1.80 ≤ z ≤ 0.20​) = P(z < 0.20) - P(z < -1.80) = 0.57926 - 0.03593 = 0.54333